The speed of a boat in still water is 10mph . it travels 15 miles upstream and 15 miles downstream in a total time of 4 hours. what is the speed of the current?
Answer by richwmiller(17219)

(10+c)*x=15
(10-c)*y=15
x+y=4
c = 5, x = 1, y = 3